4 mol NH3 + 5 mol oxygen gas-> 4 mol nitrogen monoxide+ 6 mol water if 172.68 grams of nitrogen monoxide are produced how many grams of water will also be produced


1
Expert's answer
2021-04-01T02:48:02-0400

2C8H18+25O2-->16CO2+18H2O


moles C8H18 used"= 25.1g ×\frac{1 mole}{114 g} = 0.220moles"


Assuming complete combustion, you can now figure out how many moles of CO2 are produced, since the equation tells you that 2 moles C8H18 produces 16 moles CO2.

 

moles CO2 produced "=0.220 \u00d716"


moles CO2/2 moles C8H18 "=\frac{0.220 ×16}{2}=1.76moles" of CO2 are produced